Chelsea invests a total of $17,500 in three products. She invests one part in a mutual fund which has an annual return of 11%. She invests the second part in government bonds at 7% per year. The third part she puts in CDs at 5% per year. She invests twice as much in the mutual fund as in the CDs. In the first year Chelseas investments bring a total return of $1495. How much did she invest in each product? Show work.
Answer #1

x amount invested in mutualfund......gives 11%......= 0.11x

y amount invested in goverment bond .......gives 7%.......0.07

z amount invested in in CD .......gives 5%......0.05z

.
total interest is = 1495

0.11x + 0.07y + 0.05z = 1495

11x + 7x + 5x = 149500.......(1)

.
.
total money x + y + z = 17500.......(2)

.
.
she invest twice as much in the mutual fund as in the CD

x = 2z

x - 2z = 0...........(3)

for above three equations augmented matrix is

11 7 5 149500
1 1 1 17500
1 0 -2 0

convert into  Reduced Row Eschelon Form...

Divide row1 by 11

1 7/11 5/11 149500/11
1 1 1 17500
1 0 -2 0


Add (-1 * row1) to row2

1 7/11 5/11 149500/11
0 4/11 6/11 43000/11
1 0 -2 0


Add (-1 * row1) to row3

1 7/11 5/11 149500/11
0 4/11 6/11 43000/11
0 -7/11 -27/11 -149500/11


Divide row2 by 4/11

1 7/11 5/11 149500/11
0 1 3/2 10750
0 -7/11 -27/11 -149500/11


Add (7/11 * row2) to row3

1 7/11 5/11 149500/11
0 1 3/2 10750
0 0 -3/2 -6750


Divide row3 by -3/2

1 7/11 5/11 149500/11
0 1 3/2 10750
0 0 1 4500


Add (-3/2 * row3) to row2

1 7/11 5/11 149500/11
0 1 0 4000
0 0 1 4500


Add (-5/11 * row3) to row1

1 7/11 0 127000/11
0 1 0 4000
0 0 1 4500


Add (-7/11 * row2) to row1

1 0 0 9000
0 1 0 4000
0 0 1 4500

x = 9000 amount invested in mutual fund

y = 4000  amount invested in goverment bond

z = 4500  amount invested in in CD